7. Solve the system of equations: ii) X - 3 Y = 2 X - Y = -2 Y= x - 1 X + Y = 11 iii) 2 X. +3Y=2 X - 2Y = 8
Answer #1

- x-3y= 2 0 x -y=-2 subraching from *-34 = 2 * - Y = -2 &1 (+) (-) @ ① -3y+Y = 242 -2y = 4 1 Y=4 Y = -2 considering x-3y=2 eq

(10) Y-x-1 -> x+y=11 considering o y-x-1 Xeyal -> New *+Y-11-70 tita -24--10 S 4-5 considering x + y = 11 x = 11-Y 211-5 x =

(iii) 2x+3y = 2 ->0 X-24=8 -> multiplying equation with a 2x+3y = 2 80 2x-4y = 16 >3x2 E HIG 74=-14 Y = -14 y=-2/ considering